# Zero-Waiting Load Balancing with Heterogeneous Servers in Heavy Traffic

### v1

@@ 1 Introduction
To address this question, we study a natural heterogeneous generalization of JSQ, namely the Join-the-Fastest-Shortest-Queue (JFSQ) policy as in [34, 6], in a setting with fully heterogeneous servers and in heavy-traffic regimes, including both sub and super Halfin-Whitt regimes. Our main results show that JFSQ achieves asymptotically zero waiting at steady state, with explicit system size bounds that characterize the convergence rates. The performance bounds reveal how system behavior depends on the heterogeneity level. To establish these results, we build on Stein’s method with iterative state-space peeling (ISSP) [21, 19, 20], a powerful tool for steady-state analysis of load balancing systems. However, applying it to our heterogeneous setting requires several technical innovations due to the high-dimensional and non-exchangeable state representation induced by the challenges of full heterogeneity. We directly work on the original high-dimensional state space and design a sequence of novel Lyapunov functions to show that, under JFSQ, that the servers with higher capacities are sufficiently utilized, and the steady-state system concentrates around the state space in which the total capacity of busy servers exceeds the total arrival rate. With these concentration results, we introduce a unified coupling argument that relates JFSQ to a single-server queue across both sub and super Halfin-Whitt regimes, and explicitly quantifies the delay performance. We believe these novel methodologies can be of independent interest and applied to other heterogeneous queueing systems.
@@
The performance analysis of load balancing in many-server systems has a long and rich history. For homogeneous systems, where all servers are identical, a sequence of classical load balancing policies has been proposed and analyzed, including join-the-shortest-queue (JSQ), join-the-idle-queue (JIQ) [23, 28], and power-of-$d$ (Podd) (also called JSQ-$d$) [25, 31]. The early foundational works [35, 33] proved that JSQ is delay-optimal in finite-server systems assuming Poisson arrivals and exponential service times under critical load (i.e., the arrival rate is strictly smaller than the service rate and its gap is constant independent of the system size). For the many-server asymptotic regime, a recent important work [10] analyzed JSQ in the Halfin-Whitt regime and shows that the diffusion-scaled process converges to a two-dimensional diffusion limit. The corresponding steady-state performance was analyzed in [8] via Stein’s method, and the further refined characteristics (e.g., tail probability and sensitivity) have been studied in [3, 4]. The JIQ policy has also attracted attention due to its low communication overhead [23, 28]. In [28], it was shown that JIQ achieves asymptotically vanishing delay in large-scale heterogeneous systems, though their service rates are independent of system size $N$ under the critical load regimes. For Podd, the diffusion limit of Podd is shown to converge to that of JSQ in the Halfin-Whitt regime at the process level when $d=\Theta(\sqrt{N}\log N)$ in work [26]. [21, 22] established that a class of load balancing policies, including JSQ, JIQ, and Podd with properly chosen $d$, achieve zero-waiting performance at steady state in both the sub and super Halfin-Whitt regimes. The exact asymptotic queue distribution in the super-Halfin-Whitt regime was later established in [37]. These results, however, largely rely on the assumption that all servers are homogeneous. In contrast, modern data centers and cloud platforms often exhibit substantial heterogeneity in server speeds due to differences in hardware generations and software configuration. In these heterogeneous systems, heterogeneity-unaware (or server rate-unaware) load balancing can lead to substantial inefficiencies, as jobs may be assigned to slow servers with short queues while faster servers remain underutilized. Even the classical JSQ may perform poorly [12]. This inefficiency has motivated the design of heterogeneity-aware or service-aware load balancing policies that incorporate service rates into the decision process.
@@ 2 Model and Main Results
We consider a many-server system with $N$ heterogeneous servers, where jobs arrive according to a Poisson process with rate $\lambda N$, and the service time at server $n$ is exponentially distributed with rate $1/\mu_{n}$. Without loss of generality [^1] This assumption is made for normalization. If instead $\sum_{n=1}^{N}\mu_{n}=cN$ for some constant $c>0$, we can equivalently rescale all service rates to $\mu_{n}/c$ so that the total capacity becomes $N$., the service rates satisfy $\sum_{n=1}^{N}\mu_{n}=N$. We study the many-server system under heavy-traffic regimes, where the system load approaches capacity as $\lambda=1-N^{-\alpha}$ with $\alpha\in(0,1).$ The well-known case of $\alpha=\frac{1}{2}$ corresponds to the classical Halfin-Whitt regime [14], also called quality-and-efficiency-driven (QED) regime. The regime $\alpha<\frac{1}{2}$ and $\alpha\geq\frac{1}{2}$ are named as Sub-Halfin-Whitt regime and Super-Halfin-Whitt regime, respectively. As shown in Figure 1, each server maintains a separate queue with the buffer size b1b-1 (i.e., each server can have one job in service and b1b-1 jobs in queue).
@@
The system state can be described by the vector $\mathbf{Q}(t)=(Q_{1}(t),\dots,Q_{N}(t)),$ where $Q_{i}(t)$ is the queue length of server $i$ at time $t$, including the job in service. We study the Join-the-Fastest-Shortest-Queue (JFSQ) load balancing policy, a state-aware and service rate-aware policy, where the load balancer has full knowledge of the current queue lengths $\mathbf{Q}(t)$ and the service rates $\{\mu_{n}\}$. In JFSQ, a job is routed to a server with the shortest queue length. If there are multiple such servers, the job is assigned to the one with the highest service rate. If the service rates are tied, the dispatcher selects one uniformly at random. We assume that servers are indexed in descending order of service rate, i.e., $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}\geq\epsilon$, where $\epsilon>0$ is a constant independent of $N$. Note that if $\epsilon=o(1),$ then the service time of a job on server $n$ goes to infinity as $N$ increases. We assume no such slow servers in the system, but it will become clear that this assumption can be relaxed.
Under JFSQ policy, the queue-length process $\{\mathbf{Q}(t)\}$ evolves as a continuous-time Markov chain (CTMC). We are interested in its delay performance in steady state. Let $Q_{i}$ denote the steady-state queue length at server $i$, i.e., $Q_{i}\sim Q_{i}(\infty)$. Our objective is to show if JFSQ policy can achieve “zero-waiting” performance as the system size scales and how the heterogeneity affects the delay performance. Without loss of generality, we ignore integer rounding issues and assume the existence of a critical index $N_{2}$ such that the aggregate total capacity of the fastest $N_{2}$ servers exactly supports the total arrival rate
@@
Assume $\alpha\in(0,\tfrac{1}{2})$, $r>0$, and $b\geq 2$. We further assume that the system size $N$ satisfies
@@
Assume $\alpha\in[\tfrac{1}{2},1)$, $r>0$, and $b\geq 2$. We further assume that the system size $N$ satisfies
@@
Next, we formalized the discussion above and proved the asymptotic zero waiting results. Let $\mathcal{W}$ denote the event that an incoming job is routed to a busy server, and $p_{\mathcal{W}}$ denote the probability of this event at steady-state. Let $\mathcal{B}$ denote the event that an incoming job is blocked (discarded) and $p_{\mathcal{B}}$ denote the probability of this event at steady-state. Note the ${\mathcal{B}}\subseteq{\mathcal{W}}$ because an incoming job is blocked when being routed to a busy server with $b$ jobs. Furthermore, let $W$ denote the waiting time of the jobs that are not blocked at steady-state. Denote two important parameters $k_{sub}=\frac{16(b-1)\mu_{N_{2}}}{\epsilon^{2}}$ and $k_{super}=\frac{24r(b-1)\mu_{1}}{\epsilon}$ that capture the server heterogeneity.
@@
    $\mathbb{E}[W]=\mathcal{O}\left(\frac{1}{N^{\alpha}}+\frac{k_{sub}}{N^{1-\alpha}}\right),$ $~~p_{\mathcal{W}}=\mathcal{O}\left(\frac{k_{sub}}{N^{1-\alpha}}+\frac{\log N}{\sqrt{N}}\right).$
@@
Assume $\lambda=1-N^{-\alpha}$ for $0<\alpha<1$ and buffer size b1.b-1. Given any positive integer $r$ and the service rates $\{\mu_{n}\}$ satisfy Assumptions 1 and 2. The following bound holds under JFSQ
@@
$\mathbb{E}\left[\left(\max\left\{\bar{Q}-\frac{N_{2}}{N}-\left(1-\frac{4}{k_{sub}\epsilon}\right)\frac{1}{N^{\alpha}},0\right\}\right)^{r}\right]\leq\left(\frac{k_{sub}}{N^{1-\alpha}}\right)^{r}.$ (1)
@@
Based on Theorem 2, we establish that the average queue length remains bounded under heavy-traffic regimes.
@@
    $\mathbb{E}[\bar{Q}]\leq\frac{N_{2}}{N}+\left(1-\frac{4}{k_{sub}\epsilon}\right)\frac{1}{N^{\alpha}}+\frac{k_{sub}}{N^{1-\alpha}},$
    which implies that the average queue length per server converges to $\frac{N_{2}}{N}$ as $N\to\infty$. In the homogeneous case, $N_{2}=\lambda N$, and thus $\mathbb{E}[\bar{Q}]\to\lambda$, meaning the average time spent for a job in the system approaches the normalized system load. In the heterogeneous setting, delay performance is influenced by the parameter $k_{sub}$, which depends on the degree of heterogeneity. Specifically, if $\mu_{N_{2}}/\epsilon$ is large (i.e., large $k_{sub}$ and $k_{sub}\epsilon$), the average queue length increases.
@@
    which implies that the average queue length per server converges to 1 as $N\to\infty$. Thus, all servers remain busy in the limit, and the average number of jobs waiting in the queue vanishes. As in the Sub-Halfin-Whitt regime, heterogeneity impacts queue lengths through $k_{super}$. When $\mu_{1}/\epsilon$ becomes large, $k_{super}$ also increases, leading to a larger average queue length.
3 Theretical Analysis
Recall we assume that the servers are ordered in a descending order according to their service rates, i.e., $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}.$ In our analysis, we implicitly divide the servers into multiple groups according to their aggregated service capacity. Without loss of generality, let a positive $\delta>0$ and we define $N_{1}(\delta)$ such that
@@ 3.1 Most of the fastest servers are busy
To justify the intuition “most of the fastest servers are busy” under JFSQ, we design Lyapunov functions that are directly related to the number of idle servers in two groups defined by $N_{1}$ and $N_{2}$. Intuitively, under JFSQ, the idle servers in the fastest $N_{1}$ and $N_{2}$ servers are extremely small as shown in Figure 3. The following two lemmas will further show that most idle servers among the first $N_{2}$ servers are those indexed between $N_{1}+1$ and $N_{2},$ i.e., not the fastest $N_{1}$ servers.
@@
We have for any $\delta$ satisfying $\frac{2}{N^{\alpha}}\leq\delta\leq 0.5$ such that
$\Pr\left(V(q)\leq 1+8\log^{2}N\right)$ $\geq 1-\exp\left(-{\delta}\log^{2}N\right).$
@@
We have for $\mu_{N_{2}}\leq\frac{\sqrt{N}}{\log N}$ such that
$\Pr\left(V(q)\leq 5\sqrt{N}\log N\right)\geq 1-\exp\left(-\epsilon\log^{2}N\right).$
@@ 3.2 State-space Collapse in the Heavy-traffic Regime
The first term in (4) captures the number of waiting jobs (except the jobs in service), while the second term reflects the number of busy servers (or equivalently, the complement of idle servers). The constants $A$ and $B$ are chosen based on intuition from single-server queues under heavy traffic regimes (recall $\lambda N=N-N^{1-\alpha}$ and it implies that the number of idle servers is approximated to be $\Theta(N^{1-\alpha})$ ideally).
@@
  • Super-Halfin-Whitt regime: In this regime, we expect nearly all servers to be busy. Therefore, we set $B=N$, reflecting full utilization. For $A$, we use $A=\Theta(N^{1-\alpha}\log N)$, consistent with the order of idle servers in the ideal fluid system, with an additional $\log N$ factor required for technical reasons when invoking tail probability bounds.
@@
Under JFSQ, for a large $N$ such that $N^{1-2\alpha}\geq\frac{128(r+1)(b-1)^{2}\mu_{N_{2}}\log N}{\epsilon^{2}}$ and $\epsilon\geq\frac{2(r+1)}{\log N},$ we have
$\Pr\left(V(q)\geq(1-\frac{3}{8\nu})N^{1-\alpha}\right)\geq\exp\left(-2r\log N\right).$
@@
$\Pr\left(V(q)\geq 3c_{1}N^{1-\alpha}\right)\leq\exp\left(-2r\log N\right).$

### v2

@@ 1 Introduction
To address this question, we study a natural heterogeneous generalization of JSQ, namely the Join-the-Fastest-Shortest-Queue (JFSQ) policy as in [34, 6], in a setting with fully heterogeneous servers and in heavy-traffic regimes, including both sub and super Halfin-Whitt regimes. Our main results show that JFSQ achieves asymptotically zero waiting at steady state, with explicit system size bounds that characterize the convergence rates. The performance bounds reveal how system behavior depends on the heterogeneity level. To establish these results, we build on Stein’s method with iterative state-space peeling (ISSP) [21, 19, 20], a powerful tool for steady-state analysis of load balancing systems. However, applying it to our heterogeneous setting requires several technical innovations due to the high-dimensional and non-exchangeable state representation induced by the challenges of full heterogeneity. We directly work on the original high-dimensional state space and design a sequence of novel Lyapunov functions to show that, under JFSQ, that the servers with higher capacities are sufficiently utilized, and the steady-state system concentrates around the state space in which the total capacity of busy servers exceeds the total arrival rate. With these concentration results, we introduce a unified coupling argument that relates JFSQ to the fluid model of a single-server queue across both sub and super Halfin-Whitt regimes, and explicitly quantifies the delay performance. We believe these novel methodologies can be of independent interest and applied to other heterogeneous queueing systems.
@@
The performance analysis of load balancing in many-server systems has a long and rich history. For homogeneous systems, where all servers are identical, a sequence of classical load balancing policies has been proposed and analyzed, including join-the-shortest-queue (JSQ), join-the-idle-queue (JIQ) [23, 28], and power-of-$d$ (Podd) (also called JSQ-$d$) [25, 31]. The early foundational works [35, 33] proved that JSQ is delay-optimal in finite-server systems assuming Poisson arrivals and exponential service times under critical load (i.e., the arrival rate is strictly smaller than the service rate and its gap is constant independent of the system size). For the many-server asymptotic regime, a recent important work [10] analyzed JSQ in the Halfin-Whitt regime and shows that the diffusion-scaled process converges to a two-dimensional diffusion limit. The corresponding steady-state performance was analyzed in [8] via Stein’s method, and the further refined characteristics (e.g., tail probability and sensitivity) have been studied in [3, 4]. The JIQ policy has also attracted attention due to its low communication overhead [23, 28]. In [28], it was shown that JIQ achieves asymptotically vanishing delay in large-scale heterogeneous systems, though their service rates are independent of system size $N$ under the critical load regimes. For Podd, the diffusion limit of Podd is shown to converge to that of JSQ in the Halfin-Whitt regime at the process level when $d$ is in the order of $\sqrt{N}\log N$ in work [26]. [21, 22] established that a class of load balancing policies, including JSQ, JIQ, and Podd with properly chosen $d$, achieve zero-waiting performance at steady state in both the sub and super Halfin-Whitt regimes. The exact asymptotic queue distribution in the super-Halfin-Whitt regime was later established in [37]. These results, however, largely rely on the assumption that all servers are homogeneous. In contrast, modern data centers and cloud platforms often exhibit substantial heterogeneity in server speeds due to differences in hardware generations and software configuration. In these heterogeneous systems, heterogeneity-unaware (or server rate-unaware) load balancing can lead to substantial inefficiencies, as jobs may be assigned to slow servers with short queues while faster servers remain underutilized. Even the classical JSQ may perform poorly [12]. This inefficiency has motivated the design of heterogeneity-aware or service-aware load balancing policies that incorporate service rates into the decision process.
@@ 2 Model and Main Results
We consider a many-server system with $N$ heterogeneous servers, where jobs arrive according to a Poisson process with rate $\lambda N$, and the service time at server $n$ is exponentially distributed with rate $1/\mu_{n}$. Without loss of generality [^1] This assumption is made for normalization. If instead $\sum_{n=1}^{N}\mu_{n}=cN$ for some constant $c>0$, we can equivalently rescale the arrival and service rates to $\lambda/c$ and $\mu_{n}/c,$ respectively, so that the total capacity becomes $N.$, the service rates satisfy $\sum_{n=1}^{N}\mu_{n}=N$. We study the many-server system in heavy-traffic regimes, where the system load approaches capacity as $\lambda_{N}=1-\beta N^{-\alpha}$ for any constant $\beta>0$. We let $\beta=1$ for notational simplicity and also write $\lambda$ instead of $\lambda_{N}$ when there is no ambiguity. The well-known case of $\alpha=\frac{1}{2}$ corresponds to the classical Halfin-Whitt regime [14], also called quality-and-efficiency-driven (QED) regime. The regime $\alpha<\frac{1}{2}$ and $\alpha\geq\frac{1}{2}$ are named as Sub-Halfin-Whitt regime and Super-Halfin-Whitt regime, respectively. As shown in Figure 1, each server maintains a separate queue with the buffer size b1b-1 (i.e., each server can have one job in service and b1b-1 jobs in queue).
@@
Define $\mathbb{Q}:=\left\{(q_{1},\dots,q_{N})\in\mathbb{Z}^{N}\;:\;0\leq q_{i}\leq b,\ \forall i=1,\dots,N\right\}.$ The system state can be described by the vector $\mathbf{Q}(t)=(Q_{1}(t),\dots,Q_{N}(t))\in\mathbb{Q},$ where $Q_{i}(t)$ is the queue length of server $i$ at time $t$, including the job in service. We study the Join-the-Fastest-Shortest-Queue (JFSQ) load balancing policy, a state-aware and service rate-aware policy, where the load balancer has full knowledge of the current queue lengths $\mathbf{Q}(t)$ and the service rates $\{\mu_{n}\}$. In JFSQ, a job is routed to a server with the shortest queue length. If there are multiple such servers, the job is assigned to the one with the highest service rate. If the service rates are tied, the dispatcher selects one uniformly at random. We assume that servers are indexed in descending order of service rate, i.e., $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}\geq\epsilon$, where $\epsilon>0$ is a constant independent of $N$. Note that if $\epsilon=o(1),$ then the service time of a job on server $n$ goes to infinity as $N$ increases [^2] We follow the standard big-O notation. For functions $f(N)$ and $g(N)>0$, we write $f(N)=O(g(N))$ if there exist constants $C>0$ and $N_{0}$ such that $|f(N)|\leq Cg(N)$ for all $N\geq N_{0}$; $f(N)=o(g(N))$ if $\lim_{N\to\infty}f(N)/g(N)=0$; and $f(N)=\Theta(g(N))$ if there exist constants $C_{1},C_{2}>0$ and $N_{0}$ such that $C_{1}g(N)\leq f(N)\leq C_{2}g(N)$ for all $N\geq N_{0}$.. We assume no such slow servers in the system, but it will become clear that this assumption can be relaxed.
Under JFSQ policy, the queue-length process $\{\mathbf{Q}(t)\}$ evolves as a continuous-time Markov chain (CTMC). We are interested in its delay performance in steady state. Let $Q_{i}$ denote the steady-state queue length at server $i$, i.e., $Q_{i}\sim Q_{i}(\infty)$. Our objective is to determine whether the JFSQ policy achieves “zero-waiting” performance as the system size increases, and to understand how server heterogeneity affects the resulting delay performance. For tractability, we ignore integer rounding issues and assume the existence of a critical index $N_{2}$ such that the aggregate total capacity of the fastest $N_{2}$ servers exactly supports the total arrival rate
@@
Assume $\alpha\in(0,\tfrac{1}{2})$ and $b\geq 5$. For any $r\in Z^{+},$ we assume that the system size $N$ satisfies
@@
Assume $\alpha\in[\tfrac{1}{2},1)$ and $b\geq 5$. For any $r\in Z^{+},$ we assume that the system size $N$ satisfies
@@
Next, we formalize the discussion above and proved the asymptotic zero waiting results. Let $\mathcal{W}$ denote the event that an incoming job is routed to a busy server, and $p_{\mathcal{W}}$ denote the probability of this event at steady-state. Let $\mathcal{B}$ denote the event that an incoming job is blocked (discarded) and $p_{\mathcal{B}}$ denote the probability of this event at steady-state. Note the ${\mathcal{B}}\subseteq{\mathcal{W}}$ because an incoming job is blocked when being routed to a busy server with $b$ jobs. Furthermore, let $W$ denote the waiting time of the jobs that are not blocked at steady-state. We denote two important parameters that capture server heterogeneity:
$k_{\mathrm{sub}}=\frac{16(b-1)\mu_{N_{2}}}{\epsilon^{2}}~~\textup{and}~~k_{\mathrm{super}}=\frac{24r(b-1)\mu_{1}}{\epsilon},$
which correspond to the sub-Halfin–Whitt and super-Halfin–Whitt regimes, respectively.
@@
    $\mathbb{E}[W]=\mathcal{O}\left(\frac{1}{N^{\alpha}}+\frac{k_{sub}}{N^{1-\alpha}}\right),$ $~~p_{\mathcal{W}}=\mathcal{O}\left(\frac{1}{N^{\alpha}}+\frac{k_{sub}}{N^{1-\alpha}}\right).$
@@
Assume $\lambda=1-N^{-\alpha}$ for $0<\alpha<1$ and buffer size b1.b-1. Given any positive integer $r$ and when the service rates $\{\mu_{n}\}$ satisfy Assumptions 1 and 2, the following bound holds under JFSQ
@@
$\mathbb{E}\left[\left(\max\left\{\bar{Q}-\frac{N_{2}}{N}-\left(1-\frac{4}{k_{sub}\epsilon}\right)\frac{1}{N^{\alpha}},0\right\}\right)^{r}\right]\leq{17\left(\frac{2k_{sub}r}{N^{1-\alpha}}\right)^{r}.}$ (1)
@@
Based on Theorem 2, by setting $r=1,$ we establish the upper bound of the average queue length under heavy-traffic regimes.
@@
    $\mathbb{E}[\bar{Q}]\leq\frac{N_{2}}{N}+\left(1-\frac{4}{k_{sub}\epsilon}\right)\frac{1}{N^{\alpha}}+\frac{34k_{sub}}{N^{1-\alpha}},$
    which implies that the average queue length per server is upper bounded by a quantity that scales as $\frac{N_{2}}{N}$ as $N\to\infty$. In the homogeneous case, where $N_{2}=\lambda N$, this yields $\mathbb{E}[\bar{Q}]\leq\lambda+o(1),$ indicating that the average time a job spends in the system is asymptotically controlled by the normalized system load. In the heterogeneous setting, delay performance is influenced by the parameter $k_{sub}$, which depends on the degree of heterogeneity. Specifically, if $\mu_{N_{2}}/\epsilon$ is large (i.e., large $k_{sub}$ and $k_{sub}\epsilon$), the average queue length becomes large.
@@
    which implies that the average queue length per server is upper bounded 1 as $N\to\infty$. Thus, all servers remain busy in the limit, and the average number of jobs waiting in the queue vanishes. As in the Sub-Halfin-Whitt regime, heterogeneity impacts queue lengths through $k_{super}$. When $\mu_{1}/\epsilon$ becomes large, $k_{super}$ also increases, leading to a larger average queue length.
3 Theoretical Analysis
Recall we assume that the servers are ordered in a descending order according to their service rates, i.e., $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}.$ In our analysis, we implicitly divide the servers into multiple groups according to their aggregated service capacity. Without loss of generality, let $\delta>0$ and we define $N_{1}(\delta)$ such that
@@ 3.1 Most of the fastest servers are busy
To justify the intuition that “most of the fastest servers are busy” under JFSQ, we design Lyapunov functions that are directly related to the number of idle servers in two groups defined by $N_{1}$ and $N_{2}$. Intuitively, under JFSQ, the idle servers in the fastest $N_{1}$ and $N_{2}$ servers are extremely small as shown in Figure 3. The following two lemmas will further show that most idle servers among the first $N_{2}$ servers are those indexed between $N_{1}+1$ and $N_{2},$ i.e., not the fastest $N_{1}$ servers.
@@
For any $\delta$ satisfying $\frac{2}{N^{\alpha}}\leq\delta\leq 0.5,$ we have
$\mathbb{P}\left(V(Q)\leq 1+8\log^{2}N\right)$ $\geq 1-\exp\left(-{\delta}\log^{2}N\right).$
@@
For $\mu_{N_{2}}\leq\frac{\sqrt{N}}{\log N},$ we have
$\mathbb{P}\left(V(Q)\leq 5\sqrt{N}\log N\right)\geq 1-\exp\left(-\epsilon\log^{2}N\right).$
@@ 3.2 State-space Collapse in the Heavy-traffic Regime
We remark that this Lyapunov function in (4) may take negative values, depending on the choices of $A$ and $B,$ but it is lower bounded, While we can add a sufficiently large constant to the function to make it, it becomes less intuitive. Therefore, we keep $V(\cdot)$ in the current form to better reflect the underlying intuition and will comment that the tail bound in [32] can be directly applied to a lower bounded Lyapunov function. The first term in (4) captures the number of waiting jobs (except the jobs in service), while the second term reflects the number of busy servers (or equivalently, the complement of idle servers). The constants $A$ and $B$ are chosen based on intuition from single-server queues under heavy traffic regimes (recall $\lambda N=N-N^{1-\alpha}$ and it implies that the number of idle servers is approximated to be $\Theta(N^{1-\alpha})$ ideally).
@@
  • Super-Halfin-Whitt regime: In this regime, we expect nearly all servers to be busy. Therefore, we set $B=N$, reflecting full utilization. For $A$, we use $A=\Theta(N^{1}\log N)$, consistent with the order of idle servers in the ideal fluid system, with an additional $\log N$ factor required for technical reasons when invoking tail probability bounds.
@@
Under JFSQ, for a large $N$ such that $N^{1-2\alpha}\geq\frac{128(r+1)(b-1)^{2}\mu_{N_{2}}\log N}{\epsilon^{2}}$ and $\epsilon\geq\frac{2(r+1)}{\log N}$ in Assumption 1, we have
$\mathbb{P}\left(V(Q)\geq(1-\frac{3}{8\nu})N^{1-\alpha}\right)\leq\exp\left(-2r\log N\right).$
@@
$\mathbb{P}\left(V(Q)\geq 3c_{1}N^{1-\alpha}\right)\leq\exp\left(-2r\log N\right).$